# Understanding and Improving Information Preservation in Prompt Compression for LLMs

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have enabled their successful application to a broad range of tasks. However, in information-intensive tasks, the prompt length can grow fast, leading to increased computational requirements, performance degradation, and induced biases from irrelevant or redundant information. Recently, various prompt compression techniques have been introduced to optimize the trade-off between reducing input length and retaining performance. We propose a holistic evaluation framework that allows for in-depth analysis of prompt compression methods. We focus on three key aspects, besides compression ratio: (i) downstream task performance, (ii) grounding in the input context, and (iii) information preservation. Through this framework, we investigate state-of-the-art soft and hard compression methods, showing that they struggle to preserve key details from the original prompt, limiting their performance on complex tasks. We demonstrate that modifying soft prompting methods to control better the granularity of the compressed information can significantly improve their effectiveness – up to +23% in downstream task performance, more than +8 BERTScore points in grounding, and 2.7x more entities preserved in compression.
@@
Momchil Hardalov  Laura Aina  Neha Anna John  Hang Su AWS AI Labs {momchilh, eailaura, nehajohn, shawnsu}@amazon.com
@@ 1 Introduction
A limitation of the work done on prompt compression is that it focuses mostly on downstream performance on tasks like question answering (QA, Mu et al. (2023)) or reading comprehension (RC, Li et al. (2024b)). We argue that downstream performance itself is insufficient to evaluate the quality and limitations of a certain compression method and to understand its ability to preserve information from the original context. To address this gap, we propose a holistic evaluation framework to analyze the performance of prompt compression approaches, which considers three key dimensions, beyond compression rate: (i) downstream task performance, (ii) grounding of the compressed context to the original text, and (iii) type and amount of information preserved during compression. The framework includes the evaluation on various text generation tasks with different complexity and context lengths (from 850 to 6.5K tokens), including multi-hop reasoning, conversation QA, long document summarization, and mathematical reasoning.
We use our newly proposed framework to analyze a representative hard-prompting method, LLMLingua (Jiang et al., 2023b), and a soft-prompting approach, xRAG (Cheng et al., 2024). Both methods are limited in handling long-context scenarios, exhibiting performance drops across various generation tasks and an increased number of ungrounded responses. Additionally, we demonstrate that the target LLM struggles to accurately reconstruct the original context from soft prompts, often hallucinating information and failing to regenerate crucial details. Based on these observations, we explore modifications of the xRAG soft compression method, where we incorporate fine-grained data samples during model pre-training and operate on more granular data in the generation process. Our results indicate that this approach significantly improves performance on downstream tasks (23% relative improvement), response grounding (up to 8 BERTScore F1 points), and the model’s ability to preserve information after compression (2.7 times more entities preserved).
In summary, this paper provides the following contributions: (i) it reveals major limitations of state-of-the-art soft and hard prompt compression methods; (ii) it advocates for a holistic framework to evaluate prompt compression methods beyond down-stream task performance; (iii) it presents promising directions to address the limitations of soft prompting methods, showing that manipulating the granularity of the compressed content is key to improving performance.
@@ 2 Related Work
methods compress a given text into a continuous representation in the token embedding space, referred to as special tokens (Lester et al., 2021). Approaches like AutoCompressors recursively generate summary vectors as soft prompts (Chevalier et al., 2023), while Gisting condenses prompts into transformer activations using virtual gist tokens (Mu et al., 2023). ICAE (Ge et al., 2024) and 500xCompressor (Li et al., 2024b) encode long contexts into compact memory slots, with ICAE leveraging a LoRA-adapted encoder and 500xCompressor using key-value representations for richer information retention. xRAG (Cheng et al., 2024) projects dense retrieval embeddings into the LLM’s representation space via a trainable modality bridge. It employs a frozen embedding model as the encoder, with a lightweight adapter between the encoder and the decoder LLM as the only trainable component. Unlike other soft prompting methods, xRAG keeps both the encoder and decoder frozen, making it a modular and efficient solution for rapid development. While soft prompts can achieve higher compression rates (i.e., fewer input tokens for the LLMs), they fall short in explainability, as the content encoded in the soft prompt is not directly interpretable by humans.
Most prompt compression methods are evaluated on downstream tasks, including classification (Yin et al., 2023; Chevalier et al., 2023), natural language inference (Yao et al., 2024), reasoning (Chen et al., 2023), QA (Jiang et al., 2024; Cheng et al., 2024; Xu et al., 2024), and summarization (Pan et al., 2024a; Fei et al., 2024). However, downstream performance alone does not reveal a method’s limitations or assess information loss, a key issue in prompt compression. Reconstructing the original text from the compressed prompt has been proposed as a way to measure information preservation (Li, 2023; Wingate et al., 2022; Wang et al., 2024c), but standardized metrics for this evaluation are still lacking. In this work, we go beyond downstream performance by evaluating: (i) the grounding of the generated responses in the original input, and (ii) LLM’s ability to reconstruct the input context from its compressed version. This approach enables us to study the quantity and nature of information preserved by compression methods, deriving insights into specific areas of limitations and directions for improvement.
@@ 3.1 Benchmarking Data Design
Currently, the main benchmarking of prompt compression methods (Jiang et al., 2024) focuses on a limited range of tasks such as text classification (e.g., MNLI (Yao et al., 2024); classification tasks from SuperGLUE (Lester et al., 2021; Chevalier et al., 2023)), or extractive QA/RC (e.g., DROP, Gupta et al. 2023; RACE, (Li et al., 2024b)). Solving these tasks is often possible by only capturing a high-level topic representation or the main entities. We argue that a more comprehensive evaluation is needed to truly cover the properties of the compression methods. Therefore, in our framework, we adopt generation tasks containing much longer contexts, such as long-form question answering based on a single or multiple documents, long document summarization, and conversational search, with retrieval-augmented generation (RAG) based on multiple source documents. We summarize the characteristics of the target tasks in Table 1. The proposed list is not exhaustive and should be updated with time to reflect the development of the compression methods and the LLMs.111 See Fig. 3 in Appendix C.2 for context length distribution.
@@ 3.2 Compression Quality Evaluation
Grounding is another important dimension of compression quality, providing insights into the methods’ ability to preserve key information from the context to be used in the generated responses (Kim et al., 2024). We explored different approaches to automatically evaluate grounding, to find a generalizable metric across different long-context tasks. We selected FABLES (Kim et al., 2024) because it produces results better aligned with human evaluation (see the discussion in Appendix B.2). FABLES first extracts a set of de-contextualized claims from the generated response and then rates the faithfulness of each claim given the evidence. Both steps are performed by prompting an LLM (in this work Claude 3 Haiku). For each claim, we derive a score by comparing each 10-sentence chunk from the context against the claim and then taking the maximum score across chunks. The final grounding score is the average score across all claims in the response.
@@ 4 Prompt Compression Methods Analysis
For our experiments, we select two representative soft and hard prompting methods (one for each type). For a fair comparison, we use the same target LLM model (Mistral-7B Instruct v0.2Jiang et al. 2023a) across all methods, including a setup without compression, treated as an upper bound.
@@
We adopt xRAG (Cheng et al., 2024) as our soft prompting baseline. This method was shown to work in relatively long input context scenarios, making it the most promising starting point for our research. xRAG encodes the input text into a single token passed to the target LLM. This representation is obtained from an encoder model (usually, a dense retriever), followed by a modality bridge that maps the encoder’s representations into the target LLM’s embedding space. The modality bridge is the only trainable component.
@@ 4.1 Downstream Task Performance
On this set of complex tasks, all compression methods result in sizeable performance losses (3% to 55% relative difference). The highest performance drop is observed on HotpotQA, as it requires multi-hop reasoning and aggregating relevant information. This indicates that applying compression removes crucial pieces of information needed to derive the correct answer to the questions. If we simplify the task, by retaining only the relevant paragraphs instead of the whole context (HotpotQA*), we see an improvement of 8 points for xRAG and 39 for LLMLingua, suggesting that they are not able to handle well long and noisy contexts.
On the long document summarization tasks (arXiv-sum.), we see smaller differences between the base Mistral model and the compression methods, with only a 4% relative difference in F1 BERTScore. We attribute this smaller change to the nature of the task, i.e., summarizing very long texts (~5800 words) into concise abstracts (~200 words) requires focusing on high-level information rather than details, as most of the original content is excluded from the summary.
@@
To assess the impact of the compression input granularity, we experiment with encoding the input on context-, paragraph-, and sentence-level using xRAG tokens. Results presented in Table 3 indicate that compressing context into smaller segments does not improve performance, likely because the tokens capture high-level topics without key details. This suggests that enhancing the model to handle finer details is a worthwhile direction to explore.
xRAG tokens HotpotQA HotpotQA* TriviaQA
1 per context 0.297 0.374 0.691
@@
Table 3: Exact match scores for xRAG responses with varying compression granularity.
@@ 4.2 Grounding
We evaluate how prompt compression affects the faithfulness of LLM-generated text. Depending on the task, we have different grounding texts: source documents for summarization, and background sections for QA and RC. We compare grounding scores for responses with and without compression to assess its impact. Results in Table 4 show that compression leads to a drop of 30 points in groundedness score (FABLES) on HotpotQA, and around 50 points on QuAC and arXiv-summarization. This indicates that higher compression rates used in soft prompting result in generating text that is less faithful to the context. The hallucinated content appears in the entire response/summary, including the first claim of the generated text, which intuitively contains the most important information (see Table 14 in Appendix E.2). As expected, responses generated with LLMLingua are less prone to hallucinations as the compressed input retains direct information from the original context.
@@
Table 4: FABLES grounding scores for responses generated with different methods, averaged over 5 random sets of 100 samples (stdev ($\sigma$) $<0.04$).
@@ 4.3 Information Preservation
We reconstruct the original text from the compressed xRAG representation by prompting the target LLM to recreate the information encoded in the tokens. We use a subset of xRAG pre-training prompts (Table 20 in Appendix F), and we apply them in two scenarios: (i) encoding the entire context in one xRAG token, and (ii) encoding each context sentence into separate tokens. We hypothesize that by compressing entire contexts into a single token we risk a significant information loss for longer, more complex inputs. Our evaluation of the information preservation covers only xRAG and not LLMLingua, as hard-prompting methods do not add any new content and are fully interpretable.
xRAG Sample Lengths Preserved
Data Tokens All 1-sent 5-sents 10-sents Entities
Unseen 1 per sample 0.66 0.57 0.72 0.70 0.28
@@
We evaluate our models on two datasets: “seen” data–random samples from xRAG’s pre-training set, and “unseen” data–samples from HotpotQA, QuAC, and TriviaQA. Each dataset contains 450 examples, split into three 150-example sets with 1, 5, and 10 sentences per example, respectively. Our primary metric is BERTScore, which can identify meaning-preserving paraphrases as accurate reconstructions. Table 5 indicates that the xRAG method is far from being able to reconstruct the original context (average BERTScore F1 is 0.66). On average the performance on seen and unseen examples is similar, suggesting that xRAG is learning a general representation rather than memorizing its pre-training data. Additionally, the target LLM is not able to handle more than one xRAG token, and the reconstruction scores drop by 20-30 BERTScore F1 points when using one token per sentence. Higher scores for examples containing multiple sentences likely come from the fact that xRAG pre-training data contains longer samples and the model is not able to preserve more granular pieces of information (see Figure 2 in Appendix C.1). The qualitative analysis shows that xRAG tokens primarily capture the general topic of the compressed content (see Table 23 in Appendix E.3) but fail to retain key details. This pattern is further supported by our entity preservation experiments, where we measure the fraction of text entities from the input retained in the reconstructed version (see Table 5). [^4] Table 19 in Appendix E.3 presents results per entity type.
5 Improving Soft Prompt Compression

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have enabled their successful application to a broad range of tasks. However, in information-intensive tasks, the prompt length can grow fast, leading to increased computational requirements, performance degradation, and induced biases from irrelevant or redundant information. Recently, various prompt compression techniques have been introduced to optimize the trade-off between reducing input length and retaining performance. We propose a holistic evaluation framework that allows for in-depth analysis of prompt compression methods. We focus on three key aspects, besides compression ratio: (i) downstream task performance, (ii) grounding in the input context, and (iii) information preservation. Using our framework, we analyze state-of-the-art soft and hard compression methods and show that some fail to preserve key details from the original prompt, limiting performance on complex tasks. By identifying these limitations, we are able to improve one soft prompting method by controlling compression granularity, achieving up to +23% in downstream performance, +8 BERTScore points in grounding, and 2.7× more entities preserved in compression. Ultimately, we find that the best effectiveness/compression rate trade-off is achieved with soft prompting combined with sequence-level training.
@@
Momchil Hardalov   Laura Aina   Neha Anna John   Hang Su Amazon {momchilh, eailaura, nehajohn, shawnsu}@amazon.com
@@ 1 Introduction
A limitation of the work done on prompt compression is that it focuses mostly on downstream performance on tasks like question answering (QA, Mu et al. (2023)) or reading comprehension (RC, Li et al. (2024b)). We argue that downstream performance itself is insufficient to evaluate the quality and limitations of a compression method, and to understand its ability to preserve information from the original context. To address this gap, we propose a holistic evaluation framework to analyze the performance of prompt compression approaches, which considers three key dimensions, beyond compression rate: (i) downstream task performance, (ii) grounding of the compressed context to the original text, and (iii) type and amount of information preserved during compression. The framework includes the evaluation on various text generation tasks with different complexity and context lengths (from 850 to 6.5K tokens), including multi-hop reasoning, conversation QA, long document summarization, and mathematical reasoning.
We use our newly proposed framework to analyze a representative hard-prompting method, LLMLingua (Jiang et al., 2023b), and two soft-prompting approaches, xRAG (Cheng et al., 2024) and PISCO (Louis et al., 2025). All these methods are limited in handling long-context scenarios, exhibiting performance drops across various generation tasks and an increased number of ungrounded responses. We also demonstrate that the target LLM struggles to accurately reconstruct the original context from soft prompts, particularly for xRAG, often hallucinating information and failing to regenerate crucial details. Based on these observations, we explore modifications of xRAG, where we incorporate fine-grained data samples during pre-training and operate on more granular data at generation. Our results show substantial gains: 23% improvement on downstream tasks, up to 8 BERTScore F1 points in response grounding, and 2.7× more entities preserved after compression. Still, we see limitations in the ability to integrate information across separate units. Our results show that this gap can be filled by soft prompting methods such as PISCO, which achieve the best trade-off between compression rate and effectiveness.
In summary, this paper provides the following contributions: (i) it reveals major limitations of state-of-the-art soft and hard prompt compression methods; (ii) it advocates for a holistic framework to evaluate prompt compression methods beyond downstream task performance; (iii) it presents promising directions to address the limitations of soft prompting methods, showing that manipulating the granularity of the compressed content is key to improving performance.
@@ 2 Related Work
methods compress a given text into continuous representations (Lester et al., 2021). Soft prompts can achieve higher compression rates (i.e., fewer input tokens for the LLMs), but fall short in explainability–the content encoded in the soft prompt is not directly interpretable by humans. Approaches like AutoCompressors recursively generate summary vectors  (Chevalier et al., 2023), while Gisting condenses prompts into transformer activations using virtual gist tokens (Mu et al., 2023). ICAE (Ge et al., 2024) and 500xCompressor (Li et al., 2024b) encode long contexts into compact memory slots: ICAE leverages a LoRA-adapted encoder and 500xCompressor uses key-value representations for richer information retention. xRAG (Cheng et al., 2024) projects dense retrieval embeddings into the LLM’s representation space via a trainable modality bridge. It employs a frozen embedding model as the encoder, with a lightweight adapter between the encoder and the decoder LLM as the only trainable component. Unlike other soft prompting methods, xRAG keeps both the encoder and decoder frozen, making it a modular and efficient solution for rapid development. PISCO (Louis et al., 2025) is another compressor-decoder model relying on knowledge distillation from document-based questions. While xRAG minimizes the Kullback-Leibler divergence between the logits of the encoder and target LLM, PISCO is trained using sequence-level knowledge distillation from a teacher generator.
Most prompt compression methods are evaluated on downstream tasks, including classification (Yin et al., 2023; Chevalier et al., 2023), natural language inference (Yao et al., 2024), reasoning (Chen et al., 2023), QA (Jiang et al., 2024; Cheng et al., 2024; Xu et al., 2024), and summarization (Pan et al., 2024a; Fei et al., 2024). However, downstream performance alone does not reveal a method’s limitations or assess information loss, a key issue in prompt compression. Reconstructing the original text from the compressed prompt has been proposed as a way to measure information preservation (Li, 2023; Wingate et al., 2022; Wang et al., 2024c), but standardized metrics for this evaluation are still lacking. This work go beyond downstream performance by evaluating: (i) the grounding of the generated responses in the original input, and (ii) LLM’s ability to reconstruct the input context from its compressed version. This approach enables us to study the quantity and nature of information preserved by compression methods.
@@ 3.1 Benchmarking Data Design
Currently, the main benchmarking of prompt compression methods (Jiang et al., 2024) focuses on a limited range of tasks such as text classification (e.g., MNLI (Yao et al., 2024); SuperGLUE (Lester et al., 2021; Chevalier et al., 2023)), or extractive QA/RC (e.g., DROP, Gupta et al. 2023; RACE, Li et al. 2024b). Solving these tasks is often possible by only capturing a high-level topic representation or the main entities. We argue that a more comprehensive evaluation is needed to truly cover the properties of the compression methods. Therefore, we adopt generation tasks containing much longer contexts, such as long-form QA based on multiple documents, long document summarization, and conversational search, with retrieval-augmented generation (RAG) based on multiple source documents. Table 1 summarizes the target tasks characteristics. The list is not exhaustive and should evolve with advances in compression methods and LLMs. [^1] See Fig. 3 in Appendix C.2 for context length distribution.
@@ 3.2 Compression Quality Evaluation
Grounding is another important dimension of compression quality, providing insights into a method’s ability to preserve key information from the context to be used in the generated responses (Kim et al., 2024). We explored different approaches to automatically evaluate grounding, to find a generalizable metric across different long-context tasks. We selected FABLES (Kim et al., 2024) as it produced results better aligned with human evaluation (see the discussion in Appendix B.2). FABLES first extracts a set of de-contextualized claims from the generated response and then rates the faithfulness of each claim given the evidence. Both steps are performed by prompting an LLM (Claude 3 Haiku). For each claim, we derive a score by comparing each 10-sentence context chunk against the claim and then taking the maximum score across chunks. The final score is the average across all claims in the response.
@@ 4 Prompt Compression Methods Analysis
For our experiments, we select two soft prompting and one hard prompting method, using the same target LLM – Mistral-7B Instruct v0.2 (Jiang et al., 2023a) – across all setups, with the no-compression case as an upper bound.
@@
We adopt xRAG (Cheng et al., 2024) and PISCO (Louis et al., 2025) as our soft prompting baselines. Both methods were shown to work in long context scenarios, making them the most promising starting point for our research. The input text in xRAG is encoded as a single token passed to the target LLM. This representation is obtained from an encoder model, followed by a modality bridge – the only trainable component – mapping the encoder’s representation to the target LLM’s embedding space. PISCO is a successor of xRAG containing two adapters around a backbone LLM: 1) an encoder adapter trained to compress input contexts into a set of embedding vectors, and 2) a decoder adapter providing an answer based on document embedding vectors and a query.
@@ 4.1 Downstream Task Performance
On this set of complex tasks, all compression methods result in sizeable performance losses (3% to 55% relative difference). The highest performance drop is observed on HotpotQA, as it requires multi-hop reasoning and information aggregating. This indicates that applying compression removes crucial pieces of information needed to derive the correct answer. If we simplify the task, by retaining only the relevant paragraphs instead of the whole context (HotpotQA*), we see an improvement of 8 points for xRAG and 39 for LLMLingua, indicating their difficulty with handling long, noisy contexts.
On the long document summarization tasks (arXiv-sum.), differences between the base Mistral and the compression methods are smaller, with only a 4% relative difference in F1 BERTScore. We attribute this to the nature of the task, i.e., summarizing very long texts (~5800 words) into concise abstracts (~200 words) requires focusing on high-level information rather than details.
@@
To assess the impact of the compression input granularity, we encode the input on context-, paragraph-, and sentence-level using xRAG tokens. We also run similar experiments with PISCO using separate sets of 8 embedding vectors per unit. Results presented in Table 3 indicate that compressing context into smaller segments does not improve performance for xRAG, likely because the tokens capture high-level topics without key details. However, PISCO trained with sequence-level knowledge distillation objective shows significant improvements on HotpotQA that uses longer contexts compared to the remaining two datasets when operating on smaller granularity. This suggests that enhancing the model to handle finer details is a worthwhile direction to explore for soft compression methods.
Method Encodings HotpotQA HotpotQA* TriviaQA
xRAG 1 per context 0.297 0.374 0.691
@@
PISCO 8 per context 0.318 0.589 0.738
8 per paragraph 0.512 (+61%) 0.625 (+6%)
8 per sentence 0.399 (+25%) 0.559 (-5%) 0.719 (-3%)
Table 3: Exact match scores for xRAG and PISCO responses with varying compression granularity.
@@ 4.2 Grounding
We evaluate how prompt compression affects the faithfulness of LLM-generated text. Depending on the task, we have different grounding texts: source documents for summarization, and background sections for QA and RC. We compare grounding scores for responses with and without compression to assess its impact. Results in Table 4 show that compression leads to a 30 points drop in groundedness score on HotpotQA, and around 50 points on QuAC and arXiv-summarization. This indicates that higher compression rates used in xRAG result in generating text that is less faithful to the context. The hallucinated content appears in the entire response/summary, including the first claim of the generated text, which intuitively contains the most important information (see Table 16 in Appendix E.2). As expected, responses generated with LLMLingua are less prone to hallucinations as the compressed input retains direct information from the original context. PISCO produces the most faithful responses to the source despite high compression rates, likely due to the use of sequence-level knowledge distillation from a teacher model.
@@
Table 4: FABLES grounding scores for responses generated with different methods, averaged over 5 random sets of 100 samples (stdev ($\sigma$) <0.04<0.04).
@@ 4.3 Information Preservation
We reconstruct the original text from the compressed xRAG representation by prompting the target LLM to recreate the information encoded in the tokens. Using a subset of xRAG pre-training prompts (Table 21 in Appendix F), we apply two scenarios: (i) encoding the entire context in one xRAG token, and (ii) encoding each context sentence into separate tokens. Similarly, for PISCO, we prompt the decoder to reconstruct the provided information, encoded on sample- or sentence-level, restoring as many details as possible. We hypothesize that compressing entire contexts into a single token risks significant information loss for longer, more complex inputs. Our evaluation of information preservation covers only soft prompting methods, not LLMLingua, as hard-prompting methods do not add new content and are fully interpretable.
Sample Lengths Preserved
Method Data Encodings All 1-sent 5-sents 10-sents Entities
xRAG Unseen 1 per sample 0.66 0.57 0.72 0.70 0.28
@@
Model Data Encodings Avg. PERSON GPE DATE CARD. ORG
xRAG Unseen 1 per sample 0.28 0.31 0.39 0.22 0.26 0.32
1 per sentence 0.19 0.21 0.25 0.14 0.13 0.24
Seen 1 per sample 0.25 0.16 0.38 0.24 0.38 0.21
1 per sentence 0.12 0.06 0.18 0.11 0.19 0.14
PISCO Unseen 1 per sample 0.49 0.56 0.58 0.39 0.42 0.50
1 per sent. 0.60 0.65 0.71 0.51 0.49 0.63
Table 7: Fraction of entities preserved in the reconstructed context using xRAG and PISCO models.
@@
We evaluate xRAG on two datasets: “seen” data–random samples from xRAG’s pre-training set, and “unseen” data–samples from HotpotQA, QuAC, and TriviaQA. Each dataset contains 450 examples, split into three 150-example sets with 1, 5, and 10 sentences per example, respectively. Our primary metric is BERTScore, which can identify meaning-preserving paraphrases as accurate reconstructions. Table 5 indicates that the xRAG method is far from being able to reconstruct the original context (average BERTScore F1 is 0.66). On average the performance on seen and unseen examples is similar, suggesting that xRAG is learning a general representation rather than memorizing its pre-training data. Additionally, the target LLM is not able to handle more than one xRAG token, and the reconstruction scores drop by 20-30 BERTScore F1 points when using one token per sentence.
Higher scores for examples containing multiple sentences likely come from the fact that xRAG pre-training data contains longer samples and the model is not able to preserve more granular pieces of information (see Figure 1). The qualitative analysis shows that xRAG tokens primarily capture the general topic of the compressed content (see Table 24 in Appendix E) but fail to retain key details. This pattern is further supported by our entity preservation experiments, where we measure the fraction of text entities from the input retained in the reconstructed version (last column in Table 5).
The fraction of entities’ types preserved in the reconstructed context per entity type is presented in Table 7. We observe particularly low scores for dates and numerical values. Additionally, entity preservation for people is notably poor on in-domain samples, likely due to noise or ambiguous references in the pre-training data (e.g., “Thomas Scott 1806–1816” annotated as person). In contrast, geographical locations are the most consistently preserved across all input types. Overall, our findings indicate that the model struggles with maintaining entities during prompt compression.
PISCO results are reported only for xRAG “unseen” data for fair comparison and show significantly higher information preservation—both in terms of semantic similarity between the reconstructed and original text, and the percentage of preserved entities. Notably, PISCO maintains a similar level of information retention across different encoding levels, likely due to its higher encoding capacity: while xRAG compresses information into a single token, PISCO uses 8 embeddings.
4.4 Compression Computational Efficiency
We evaluate the computational cost of various prompt compression methods by measuring the total number of floating-point operations (FLOPs) required to compress a standard input sample (context documents for a randomly selected HotpotQA dataset sample). The results, summarized in Table 8, provide a comparison of the efficiency of each compression method, highlighting trade-offs between compression strategy and computational demand. Dominating operation remains the number of matrix multiplication. The results show that all compression methods used in our experiments are comparable in terms of efficiency.
Method MFLOPs
xRAG $3.7x10^{7}$
PISCO $1.3x10^{7}$
LLMLingua $6.6x10^{6}$
Table 8: Computational cost of different prompt compression methods measured in floating-point operations (FLOPs). Values are reported in MFLOPs, representing the total number of FLOPs required for processing a single input sample.
5 Improving xRAG